Wednesday, October 16, 2019, 11:00 EDT

From Hall A Wiki
Jump to: navigation, search

Back to Main Compton Page >> Compton Meetings

*********BLUEJEANS CALL-IN INSTRUCTIONS*********
----------------------------------------------------
1. Dial Toll-Free Number: 1-888-240-2560 (U.S. & Canada)
2. International participants dial:
Toll Number: 1-408-740-7256
3. Enter meeting ID:140287970

Desktop sharing:
https://bluejeans.com/140287970 

Agenda

Meeting room = Cebaf Center F227

General

  • PSS certification in hall this week
  • Hall will be locked up sometime around Nov. 11 for high current tests
  • Spin dance is tentatively scheduled around Nov. 20

Electron detector

  • Detector installation
  • QWAD adapter boards
  • Modified flange should be ready in a couple weeks
  • Counting DAQ


Laser

  • New LOSP draft submitted all users will need Lock and Tag training
  • See email about update laser safety training. Complete by end of month.
  • Re-started laser system
  • Checked new seed laser

Photon Detector

  • PREX checks
  • CREX preparations

Analysis

  • ACC0 - laser-cycle analysis
  • ACC4 - laser cycle analysis

Minutes

  1. General
    1. PSS Certification might require people leave the hall.
    2. High current tests currently scheduled November 11, Spin dance moved to November 15.
  1. Electron Detector
    1. We should get it in ASAP, we don't want to break vacuum too close to run start since the pump down is expected to take 1 week.
    2. Robin says modified flange should take 3 weeks, but since it hasn't been ordered already we will not install it but we'll still order it.
    3. Install 3x 500 um planes and 1x 1mm thick plane.
    4. Michael shared a presentation:
      1. Modeled French electronics in a simulation. Signals are connected to HV through 1 MOhm. Peak current 0.7 uA. Adding extra capacitance to model almost destroys signal.
      2. Adding a preamplifier in the vacuum increases signal 7x.
      3. QWADS need different biasing. Bias backplane with 200 V through 1 MOhm. Signals readout straight to QWAD board. Even additional capacitance seems tolerable.
      4. Final decision: Make an adaptor board to connect French crate connector to QWAD connector. Who can manufacture these boards? Does Juliette have an unmodified QWAD to test with? We should aim to use the maximum detector bias. We should make 2 sets of board with 2 attenuation factors.
    5. Counting DAQ: FADC readout problem solved.
  1. Laser
    1. New laser safety procedures coming in soon. New walkthrough and other training required.
    2. Installed new seed laser and locking was good. Beam is no longer round entering the cavity and the cavity coupling and power are about 1/2 of what they used to be. Old seed laser was installed and it is locking well now too (but at low coupling/cavity power). Later the new-new-10W fibre amp was pulled and the new-old-10W fibre amp was installed. If the 10WNN is not activated we should send it to IPG to fix the key. elog:550
  1. Photon Detector
    1. LED studies: large noise when the LED system is on. Might be fixed by fixing cable connections. log:3734661
  1. Analysis
    1. Jumps in the accumulator are not understood. Need a realistic uncertainty for those cycles with large changes.
    2. Collimator position has been fit elog:548 and analyzing power extracted from simulation for different periods elog:549.